OpenStudy (anonymous):

Find the difference. 3x2y - x2y

OpenStudy (anonymous):

Typically, if you subtract x2y from 3x2y, you'll be left with just 3.

OpenStudy (agent0smith):

@preskill89 not quite... :/ @AllieDanielle what is 3y - y?

OpenStudy (anonymous):

Apologies on my critical error. I presume you mean: 3x^2y=x^2y Therefore, your answer would be: 2x^2y

OpenStudy (agent0smith):

I think it's just an expression, \[\Large 3x^2y - x^2y\]needs to be simplified by combining like terms.

OpenStudy (unklerhaukus):

3x^2y - x^2y = (3-1) x^2y